OpenStudy (anonymous):

A body of mass 1 kg is kept at rest. A constant force of 6.0 N starts acting on it. Find the time taken by the body to move through a distance of 12 m.

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

F=ma f=1 * a= 6n a=6 s=ut+1/2 at^2 u=0 so find s=1/2 at^2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then any hint pls tell

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

@ParthKohli help!

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

@!Developer! follow theyan

OpenStudy (anonymous):

first find acceleration

OpenStudy (lgbasallote):

what more do you need?

OpenStudy (anonymous):

and use s = ut + 1/2 a t^2

OpenStudy (anonymous):

then the answer comes 2 s as 't' then how it will come

OpenStudy (anonymous):

12 = 1/2 * 6 * t^2 12*2/6 = t^2 t= 2s

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome...))

OpenStudy (anonymous):

thanks once again

OpenStudy (anonymous):

it is 24/6= 4 s

OpenStudy (anonymous):

u see it t^2

OpenStudy (turingtest):

this is a physics question please post questions in their proper category and wait patiently for help

OpenStudy (anonymous):

t^2 = 4 t = sqrt4 = 2

OpenStudy (turingtest):

@!Developer!

OpenStudy (anonymous):

If u r satisfied plzz give medal!!

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

I got confused

OpenStudy (anonymous):

thxxxx